# What Should We Engineer in Prompts? Training Humans in Requirement-Driven LLM Use

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Applied computing Computer-assisted instructionccs: Computing methodologies Natural language generation
@@ 1. Introduction
General-purpose AIs like large language models (LLMs) have evolved from simple next-word predictors (Brown, 2020) to powerful assistants capable of fulfilling complex user needs (Qin et al., 2024; Ouyang et al., 2022). This improvement in LLM’s ability to follow instructions has encouraged users to delegate increasingly intricate tasks to these models. In the early days of prompt engineering, users primarily focused on refining the wording of simple instructions to improve LLM output quality (Jiang et al., 2020; Shin et al., 2020). Today, prompts resemble detailed “essays” that define LLM behaviors. Rather than one-off small requests, these prompts start to power the generation of complex applications, and everyday users can write complex prompts to either generate code for softwares (Wodecki, 2024; Yang et al., 2024a) or tailor general-purpose LLMs into special purpose LLM applications (Zhang et al., 2024). For instance, an LLM application (or a GPTs) like Trip Advisor  [^1] https://chatgpt.com/g/g-K38J9feSL-trip-advisor can be built solely using prompts similar to Figure 1C. These LLM applications function similarly to traditional software and are accessible through LLM App Stores like GPT Store [^2] https://openai.com/index/introducing-the-gpt-store/ and FlowGPT [^3] https://flowgpt.com/, where they can be reused with a single click. In just a few months, over 30,000 “LLM app developers” have created millions of GPT apps, with the most popular ones being used more than 5 million times (Zhang et al., 2024). This trend signals a future of end-user programming through natural language, where anyone can create reusable programs just by writing prompts (Feng et al., 2024).
@@
The rise of optimizers makes us think: What aspects of prompt engineering should humans still master? We argue that the remaining part of manual prompt creation is akin to the core step before engineering in a design process: requirement specification. A requirement is a description of user goals; it defines and documents how a system should behave, including expected inputs and outputs (Lamsweerde, 2009) (e.g., ‘‘Start the response with a tl;dr’’ for Trip Advisor in Figure 1B). Requirements are the backbone of program functionalities, and user-specified requirements in prompts (e.g., Figure 2) are the driving forces behind complex LLM adaptations. Customized requirements are hard to predict, and thus difficult to fully automate and must be specified by humans. We argue that in a future where prompt programs are common, individuals should develop proficiency in requirement articulation. In other words, we need to shift our focus towards Requirement-Oriented Prompt Engineering, a novel paradigm that we call “ROPE.
Articulating clear and complete requirements is a known challenge in fields such as product design and software engineering (Liu et al., 2023b; Nam et al., 2023; Wang and Zeng, 2009), and poor requirements often cause inefficiency, frustrations, and software failures (Zamfirescu-Pereira et al., 2023; Millett et al., 2007). This difficulty is amplified by the unpredictable nature of LLMs, which complicates controlling outputs (Arzberger et al., 2024; Chopra et al., 2023). Even experts need multiple iterations to refine requirements (Shankar et al., 2024a), and novices frequently make requirement mistakes (Babe et al., 2023). Recognizing both the importance and difficulty of this task, we pose the research question: Can we help end-users better instruct LLMs to achieve their goals through ROPE training?
@@
To answer this question, we develop the ROPE paradigm (Section 3), along with the requirement-focused training and assessments to (1) help users practice writing complete and correct requirements, and (2) deepen our understanding of how requirements affect the quality of LLM outputs. Specifically, we formalize tasks (Section 4.1) with complex requirements (e.g., game development), and design corresponding assessments (Section 4.2) that can measure the requirement quality, the LLM output quality, and their correlations. We also design and implement an interactive training mechanism (Section 4.3) that disentangles requirement articulation from other adjacent prompting aspects. Our system supports deliberate practice on requirement refinement by generating immediate feedback grounded on reference requirements.
We conduct a randomized controlled user study (Section 5) with 30 prompting novices, comparing the requirements users write before and after training, and against a control group that received standard prompt engineering training. We confirm that our requirement-focused training is effective (Section 6): novices receiving this training significantly improved their requirement performances by two-fold in the pre-to-post test, achieved significantly more gains than conventional prompt engineering training, and learned to iterate on prompts in a more structured way towards requirements. Moreover, we confirm that requirements are indeed important for the LLM generation: There is a strong positive correlation between the requirement qualities and the LLM output qualities. Importantly, the training gains are complementary to potential improvements from optimizers: Using a popular prompt optimizer, we observe that optimizers can indeed improve requirement quality, but they cannot close the gap between the baseline and the training groups. We conclude with an in-depth discussion (Section 7) on future directions for human training and optimizer design to enhance human-LLM complementarity in ROPE.
@@
  • We provide both quantitative and qualitative insights for the future of prompting, highlight the relationship between the quality of requirements and LLM performance, and discuss current optimizers’ limitations and potentials.
@@ 2.1. Current Prompt Engineering Practices and Challenges
Prompt engineering (PE) has been essential for crafting effective input instructions to guide LLMs toward generating desired outputs (Liu et al., 2021; Bach et al., 2022; Zamfirescu-Pereira et al., 2023). However, the non-deterministic nature of LLMs makes it challenging for humans to predict LLMs’ behavior across prompts (Ouyang et al., 2024; Denny et al., 2024; Sun et al., 2022). This leads to wasted time on unproductive strategies, such as trivial wording changes (Feldman and Anderson, 2024; Denny et al., 2024; Zamfirescu-Pereira et al., 2023). While some challenges can be mitigated through automation (see Section 2.2), e.g., automating word choices, human requirements (some also refer to as constitutions or principles in literature (Petridis et al., 2024)) remain crucial for customized or specialized tasks (Liu et al., 2021). We define these tasks as LLM-hard, as a simple prompt cannot produce a satisfactory response. Developing reusable prompts for customized chatbots or GPTs (Bach et al., 2022; Zamfirescu-Pereira et al., 2023) is a common LLM-hard task. While a prompt is an input to LLMs, a reusable prompt is a set of instructions for recurring goals, essentially functioning as a prompt program. As more users create reusable prompts, the ability to write good prompt programs with clear requirements becomes increasingly important.
@@
Existing PE training emphasizes the importance of explicitly stating requirements within prompts (Meskó, 2023; Schmidt et al., 2024; Santana, 2024), yet there is still a lack of focused training on requirement articulation skills for everyday users. Without training, non-experts often make mistakes such as missing requirements and writing conflicting requirements in prompts (Feldman and Anderson, 2024; Nam et al., 2023; Prather et al., 2024; Mahdavi Goloujeh et al., 2024). Even experts need many iterations to improve their requirements for complicated LLM tasks (Shankar et al., 2024a). However, training novices on PE is difficult; for instance, novice programmers may not improve at prompting within a 75 minute study when provided with test cases and generated code as feedback (Nguyen et al., 2024). We aim to address this gap by proposing a ROPE training for end-users to improve their requirements engineering skills in the context of prompt creation.
@@ 2.2. Instruction Following for Foundation Models
Optimizing LLM performance on customized tasks typically follows two approaches: improving the models directly, or refining prompts with models fixed. The former has enabled the versatility of models like GPT-4, [^5] https://openai.com/index/gpt-4/ Gemini, [^6] https://gemini.google.com/ and LLaMA 3, [^7] https://www.llama.com/ which are fine-tuned through instruction tuning (Ouyang et al., 2022) and further aligned with human preferences using Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). For the latter, the NLP community has been exploring LLMs’ capabilities to follow requirements (or “constraints”). Various datasets have been proposed to assess whether LLMs can fulfill constraints, such as IFEval (Zhou et al., 2023), INFOBench (Qin et al., 2024), and FollowBench (Jiang et al., 2024). These datasets focus on evaluating models’ abilities to handle compositional constraints, and have demonstrated promising advancements.
@@ 3. The ROPE Paradigm
We offer our definition of Requirement-Oriented Prompt Engineering (ROPE) and describe why, when, and who need it, and why we propose a training toward ROPE in this work. In short, ROPE represents a paradigm shift in how we interact with LLMs, focusing on the importance of crafting accurate and complete requirements to achieve better results, especially for complex, customized tasks.
@@
Ideally, we hope LLM users would naturally refine requirements in prompts based on unsatisfactory model outputs. In reality, this is challenging. Novices struggle to understand what constitutes a requirement in prompt engineering (Zamfirescu-Pereira et al., 2023), while even experienced users struggle at extracting and expressing requirements appropriately. For example, users find it difficult to translate raw observations on model outputs (“I don’t like how the chatbot didn’t introduce itself”) to clear requirements (“Introduce yourself at the start of the conversation, and state what you can help with”) (Petridis et al., 2024; Liu et al., 2023a); Users also cannot decide on the right level of specificity and abstraction (Liu et al., 2023b), e.g., use overly general keywords where more granular, domain-specific requirements are needed for LLM-hard tasks (Nguyen et al., 2024; Nam et al., 2023).
@@ 5. User Study Design
Our user study — for either the ROPE or the PE group — is 1.5 hours long. In the study, participants first completed a two-minute pre-survey, providing demographic information and rating their prompting experiences (e.g., confidence and self-perceived productivity using ChatGPT) with 7-level Likert Scale questions. Then, they started a pre-test (20 minutes) to write prompts to replicate key features on existing programs (Table 1) — one Game (Tic-Tac-Toe or Connect4) and one GPTs (Outline Assistant or Trip Advisor). They proceeded to the 40-minute training session (different interfaces depending on experiment group) on two tasks (Tetris and Email Proofreader), followed by a post-test that contained different Game and GPTs tasks (20 minutes).141414 The pre- and post- assessments each contained one Game task and one GPTs task (Section 4.1), counterbalanced to mitigate problem sequencing bias or different task difficulties. We found that the two versions of the tests have comparable difficulties from pre-test results (an average overall score of 21.0% and 25.6% for each version). To further capture participants’ iterative prompting behaviors, for the post-test GPTs task, participants were also asked to feed their prompt into ChatGPT and iterate their original prompt by observing the ChatGPT outputs. They completed the study with another two-minute post-survey, providing feedback with open-ended or Likert Scale questions on their experience and perceptions during the training. Participants were compensated with a \$20 Gift Card. Please refer to supplemental materials for our study materials.
@@
To understand participants’ perceived experiences, prompting behaviors, and learning outcomes, we collected participants’ prompt responses, survey answers, time on task, as well as interaction log data during the training phase. Afterward, we collected the LLM outputs by feeding participant prompts into OpenAI GPT-4o. [^16] We used a temperature of 0 to generate deterministic outputs, please refer to our supplemental materials for more details. Using the evaluation method and rubrics we developed for each of the assessment tasks Section 4.2, we graded the pre- and post- tests using users’ original prompts for all tasks from the test, on a scale of 0% to 100%. We calculated the Overall Scores as the average of the Requirement Quality and LLM Output Quality scores.
@@ 6.1. Learning Gain: Requirement-focused Training Helps Students Instruct LLMs
We start by answering our central research questions: Can we help end-users better instruct LLMs through requirement-focused training? We quantitatively measured learning gains by capturing participant performances from the pre-test to post-test within each experiment group, and evaluated the effectiveness of the two training approaches by comparing the ROPE and PE group. We further unpacked the quantitative results through analyses of participants’ Likert Scale ratings, as well as the transcribed comments during their training session and in post surveys.
@@
Using a two-tailed paired t-test, [^18] We did a linear mixed effects regression to account for the correlation across participants and task (both as random effects) of the repeated measures for the pre- and post-test times and for the requirement and LLM output scores. We found the same pattern of significant results as revealed by the t-tests. we found that the overall scores for participants in the ROPE group significantly improved by 19.0% [^19] Requirement Quality score improved by 25.4% and LLM Output Quality score improved by 12.7%, both significant ($p<0.05$). from pre- to post-test ($p<0.001$, from $21.9\%\pm 14.6\%$ to $40.9\%\pm 18.4\%$), an almost two-fold increase. This demonstrated that requirement-focused training enabled participants to instruct LLMs more effectively, and we achieved our desired goal: to concentrate users’ attention on only requirement iterations during the training, but make sure users can still apply their learning in overall prompt writing. Analyzing requirement defects in ROPE participants’ prompts, we observed a noticeable decrease in omission errors (from 5.6 to 3.2 per participant), but a slight increase in commission errors (from 0.5 to 0.7 per participant). This indicated that participants became more aware of requirements, but need additional training for requirements clarity and accuracy.
@@
These prompt also revealed learning gain differences. A two-sample t-test showed that the ROPE achieved significantly higher learning gains compared to the PE group ($p<0.001$). In fact, while ROPE significantly improved (19.0% as mentioned above), we did not observe much gain in PE group’s pre-to-post score (0.7%). The contrast suggested that novices indeed could not acquire requirement articulation skills through standard prompting training alone.

### v3

@@ Abstract.
conference: ; 2025; ACMcopyright: ccjournal: TOCHIjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3731756ccs: Human-centered computing Interactive systems and toolsccs: Applied computing Computer-assisted instructionccs: Computing methodologies Natural language generation
@@
Diagram showing the ROPE (Requirement-Oriented Prompt Engineering) framework. It includes requirement-focused training with feedback mechanisms and pre/post-training assessment tasks. The training provides feedback, summarizes requirements, and evaluates prompt output. A study result shows increased learning gains with ROPE training compared to control.
@@ 1. Introduction
General-purpose AIs like large language models (LLMs) have evolved from simple next-word predictors (Brown et al., 2020) to powerful assistants capable of fulfilling complex user needs (Qin et al., 2024; Ouyang et al., 2022). This improvement in LLM’s ability to follow instructions has encouraged users to delegate increasingly intricate tasks to these models. In the early days of prompt engineering, users primarily focused on refining the wording of simple instructions to improve LLM output quality (Jiang et al., 2020; Shin et al., 2020). Today, prompts resemble detailed “essays” that define LLM behaviors. Rather than one-off small requests, these prompts start to power the generation of complex applications, and everyday users can write complex prompts to either generate code for softwares (Wodecki, 2024; Yang et al., 2024b) or tailor general-purpose LLMs into special purpose LLM applications (Zhang et al., 2024). For instance, an LLM application (or a GPTs) like Trip Advisor  [^1] https://chatgpt.com/g/g-K38J9feSL-trip-advisor can be built solely using prompts similar to Figure 1C. These LLM applications function similarly to traditional software and are accessible through LLM App Stores like GPT Store [^2] https://openai.com/index/introducing-the-gpt-store/ and FlowGPT [^3] https://flowgpt.com/, where they can be reused with a single click. In just a few months, over 30,000 “LLM app developers” have created millions of GPT apps, with the most popular ones being used more than 5 million times (Zhang et al., 2024). This trend signals a future of end-user programming through natural language, where anyone can create reusable programs just by writing prompts (Feng et al., 2024).
@@
The rise of optimizers makes us think: What aspects of prompt engineering should humans still master? We argue that the remaining part of manual prompt creation is akin to the core step before engineering in a design process: requirement specification. A requirement is a description of user goals; it defines and documents how a system should behave, including expected inputs and outputs (Lamsweerde, 2009) (e.g., ‘‘Start the response with a tl;dr’’ for Trip Advisor in Figure 1B). Requirements are the backbone of program functionalities, and user-specified requirements in prompts (e.g., Figure 2) are the driving forces behind complex LLM adaptations. Customized requirements are hard to predict, and thus difficult to fully automate and must be specified by humans. We argue that in a future where prompt programs are common, individuals should develop proficiency in requirement articulation. In other words, we need to shift our focus towards Requirement-Oriented Prompt Engineering, an emerging paradigm that we call “ROPE.
Articulating clear and complete requirements is a known challenge in fields such as product design and software engineering (Liu et al., 2023a; Nam et al., 2024; Wang and Zeng, 2009), and poor requirements often cause inefficiency, frustrations, and software failures (Zamfirescu-Pereira et al., 2023; Millett et al., 2007). This difficulty is amplified by the unpredictable nature of LLMs, which complicates controlling outputs (Arzberger et al., 2024; Chopra et al., 2023). Even experts need multiple iterations to refine requirements (Shankar et al., 2024a), and novices frequently make requirement mistakes (Babe et al., 2024). Recognizing both the importance and difficulty of this task, we pose the research question: Can we help end-users better instruct LLMs to achieve their goals through ROPE training?
@@
[…]
Side-by-side comparison of a human-written prompt and its optimized version for a Trip Advisor task. The human prompt is informal and fragmented, whereas the optimized prompt is structured with sections. The optimized version improves fluency and structure using role-play and chain-of-thought methods.
To answer this question, we present the ROPE paradigm as a future human-LLM partnering strategy (Section 3), and we develop requirement-focused training and assessments to (1) help users practice writing complete and correct requirements, and (2) deepen our understanding of how requirements affect the quality of LLM outputs. Specifically, we formalize tasks (Section 4.1) with complex requirements (e.g., game development), and design corresponding assessments (Section 4.2) that can measure the requirement quality, the LLM output quality, and their correlations. We also design and implement an interactive training mechanism (Section 4.3) that disentangles requirement articulation from other adjacent prompting aspects. Our system supports deliberate practice on requirement refinement by generating immediate feedback grounded on reference requirements.
We conduct a randomized controlled user study (Section 5) with 30 prompting novices, comparing the prompts users write before and after training, and against a control group that received standard prompt engineering training. We confirm that our requirement-focused training is effective (Section 6): novices receiving this training significantly improved their requirement performances by two-fold in the pre-to-post test, achieved significantly more gains than conventional prompt engineering training, and learned to iterate on prompts in a more structured way towards requirements. Moreover, we confirm that requirements are indeed important for the LLM generation: There is a strong positive correlation between the requirement qualities and the LLM output qualities. Importantly, the training gains are complementary to potential improvements from optimizers and models: Using a popular prompt optimizer, we observe that optimizers can indeed improve requirement quality, but they cannot close the gap between the baseline and the training groups; Moreover, when switching from GPT-4o to the more advanced reasoning model o3-mini, the model’s improved instruction-following capabilities made its outputs more reflective of users’ explicit requirements, rather than rendering requirements obsolete. We conclude with an in-depth discussion (Section 7) on future directions for human training and optimizer design to enhance human-LLM complementarity in ROPE.
@@
  • We provide both quantitative and qualitative insights for the future of prompting, highlight the relationship between the quality of requirements and LLM performance across model advancements, and discuss current optimizers’ limitations and potentials.
@@ 2.1. Current Prompt Engineering Practices and Challenges
Prompt engineering (PE) has been essential for crafting effective input instructions to guide LLMs toward generating desired outputs (Liu et al., 2023c; Bach et al., 2022; Zamfirescu-Pereira et al., 2023). However, the non-deterministic nature of LLMs makes it challenging for humans to predict LLMs’ behavior across prompts (Ouyang et al., 2025; Denny et al., 2024; Sun et al., 2022). This leads to wasted time on unproductive strategies, such as trivial wording changes (Feldman and Anderson, 2024; Denny et al., 2024; Zamfirescu-Pereira et al., 2023). While some challenges can be mitigated through automation (see Section 2.2), e.g., automating word choices, human requirements (some also refer to as constitutions (Petridis et al., 2024) or principles (Louie et al., 2024) in literature) remain crucial for customized or specialized tasks (Liu et al., 2023c). We define these tasks as LLM-hard, as a simple prompt cannot produce a satisfactory response. Developing reusable prompts for customized chatbots or GPTs (Bach et al., 2022; Zamfirescu-Pereira et al., 2023) is a common LLM-hard task. While a prompt is an input to LLMs, a reusable prompt is a set of instructions for recurring goals, essentially functioning as a prompt program. As more users create reusable prompts, the ability to write good prompt programs with clear requirements becomes increasingly important.
@@
Existing PE training emphasizes the importance of explicitly stating requirements within prompts (Meskó, 2023; Schmidt et al., 2024; Santana, 2024), yet there is still a lack of focused training on requirement articulation skills for everyday users. Without training, non-experts often make mistakes such as missing requirements and writing conflicting requirements in prompts (Feldman and Anderson, 2024; Nam et al., 2024; Prather et al., 2024; Mahdavi Goloujeh et al., 2024). Even experts need many iterations to improve their requirements for complicated LLM tasks (Shankar et al., 2024a). However, training novices on PE is difficult; for instance, novice programmers may not improve at prompting within a 75 minute study when provided with test cases and generated code as feedback (Nguyen et al., 2024). We aim to address this gap by proposing a ROPE training for end-users to improve their requirements engineering skills in the context of prompt creation.
@@ 2.2. Instruction Following for Foundation Models
Optimizing LLM performance on customized tasks typically follows two approaches: improving the models directly, or refining prompts with models fixed. The former has enabled the versatility of models like GPT-4, [^5] https://openai.com/index/gpt-4/ Gemini, [^6] https://gemini.google.com/ and LLaMA 3, [^7] https://www.llama.com/ which are fine-tuned through instruction tuning (Ouyang et al., 2022) and further aligned with human preferences using Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). For the latter, the NLP community has been exploring LLMs’ capabilities to follow requirements (or “constraints”) with various techniques such as decomposition (least-to-most prompting (Zhou et al., 2022) and AI chains (Wu et al., 2022)) for complex tasks. Various datasets have been proposed to assess whether LLMs can fulfill constraints, such as IFEval (Zhou et al., 2023), INFOBench (Qin et al., 2024), and FollowBench (Jiang et al., 2024). These datasets focus on evaluating models’ abilities to handle compositional constraints, and have demonstrated promising advancements.
@@ 3. The ROPE Paradigm
We offer our definition of Requirement-Oriented Prompt Engineering (ROPE) and describe why, when, and who need it, and why we propose a training toward ROPE in this work. ROPE is a human-LLM partnering strategy where humans maintain agency and control of the goals via specifying requirements for LLM prompts. ROPE represents an emerging paradigm in how we interact with LLMs, focusing on the importance of crafting accurate and complete requirements to achieve better results, especially for complex, customized tasks.
@@
[…]
A conceptual diagram of the ROPE paradigm. A human refines requirements, which are passed to a compiler that optimizes prompts for large language models (LLMs). Arrows indicate the flow between human and compiler and from compiler to LLMs.
@@
Ideally, we hope LLM users would naturally refine requirements in prompts based on unsatisfactory model outputs. In reality, this is challenging. Novices struggle to understand what constitutes a requirement in prompt engineering (Zamfirescu-Pereira et al., 2023), while even experienced users struggle at extracting and expressing requirements appropriately. For example, users find it difficult to translate raw observations on model outputs (“I don’t like how the chatbot didn’t introduce itself”) to clear requirements (“Introduce yourself at the start of the conversation, and state what you can help with”) (Petridis et al., 2024; Liu et al., 2023b); Users also cannot decide on the right level of specificity and abstraction (Liu et al., 2023a), e.g., use overly general keywords where more granular, domain-specific requirements are needed for LLM-hard tasks (Nguyen et al., 2024; Nam et al., 2024).
@@ 4.2. Assessment Design: Requirement-Focused Intrinsic and Extrinsic Evaluation
[…]
Screenshots of the ROPE training interface showing feedback types: (A) Chat-based hints for requirement defects, (B) Reference requirement examples, and (C) LLM output counterfactual to reflect on incorrect or ambiguous inputs. The scenario focuses on developing Tetris game requirements.
@@ 5. User Study Design
Our user study — for either the ROPE or the PE group — is 1.5 hours long. In the study, participants first completed a two-minute pre-survey, providing demographic information and rating their prompting experiences (e.g., confidence and self-perceived productivity using ChatGPT) with 7-level Likert Scale questions. Then, they started a pre-test (20 minutes) to write prompts to replicate key features on existing programs (Table 1) — one Game (Tic-Tac-Toe or Connect4) and one GPTs (Outline Assistant or Trip Advisor). They proceeded to the 40-minute training session (different interfaces depending on experiment group) on two tasks (Tetris and Email Proofreader), followed by a post-test that contained different Game and GPTs tasks (20 minutes).141414 The pre- and post- assessments each contained one Game task and one GPTs task (Section 4.1), counterbalanced to mitigate problem sequencing bias or different task difficulties. We found that the two versions of the tests have comparable difficulties from pre-test results (an average overall score of 21.0% and 25.6% for each version). To further capture participants’ iterative prompting behaviors, for the post-test GPTs task, participants were also asked to feed their prompt into ChatGPT and iterate their original prompt by observing the ChatGPT outputs. They completed the study with another two-minute post-survey, providing feedback with open-ended or Likert Scale questions on their experience and perceptions during the training. Participants were compensated with a \$20 Amazon Gift Card. Please refer to supplemental materials for our study materials.
@@
To understand participants’ perceived experiences, prompting behaviors, and learning outcomes, we collected participants’ prompt responses, survey answers, time on task, as well as interaction log data during the training phase. Afterward, we collected the LLM outputs by feeding participant prompts into OpenAI GPT-4o. [^16] We used a temperature of 0 for GPT-4o to generate deterministic outputs, please refer to our supplemental materials for more details. Using the evaluation method and rubrics we developed for each of the assessment tasks Section 4.2, we graded the pre- and post- tests using users’ original prompts for all tasks from the test, on a scale of 0% to 100%. We calculated the Overall Scores as the average of the Requirement Quality and LLM Output Quality scores. To further understand the effect of model advancement, we collected LLM outputs with the newer, more powerful reasoning model OpenAI o3-mini [^17] https://openai.com/index/openai-o3-mini/ with the default settings. and conducted the same analysis. The results and comparisons against GPT-4o are presented in Section 6.2.
@@ 6.1. Learning Gain: Requirement-focused Training Helps Students Instruct LLMs
We start by answering our central research questions: Can we help end-users better instruct LLMs through requirement-focused training? We quantitatively measured learning gains by capturing participant performances from the pre-test to post-test within each experiment group, and evaluated the effectiveness of the two training approaches by comparing the ROPE and PE group. We further unpacked the quantitative results through analyses of participants’ Likert Scale ratings, [^19] We investigated whether the learning gains are associated with any demographic factors like age and did not find any significant indicator. There is a weak negative correlation between self-perceived familiarity with LLM and learning gains, which means that the lower LLM familiarity the participants rated, the more they learn from the training ($\rho=-0.2$). as well as the transcribed comments during their training sessions and in post surveys.
@@
Using a two-tailed paired t-test, [^20] We did a linear mixed effects regression to account for the correlation across participants and task (both as random effects) of the repeated measures for the pre- and post-test times and for the requirement and LLM output scores. We found the same pattern of significant results as revealed by the t-tests. we found that the overall scores for participants in the ROPE group significantly improved by 19.1% [^21] Requirement Quality score improved by 25.4% and LLM Output Quality score improved by 12.7%, both significant ($p<0.05$). Pre-test performances were not correlated with learning gains. from pre- to post-test ($p<0.001$, from $21.9\%\pm 14.6\%$ to $40.9\%\pm 18.4\%$), an almost two-fold increase (as shown in Figure 5a). This demonstrated that requirement-focused training enabled participants to instruct LLMs more effectively, and we achieved our desired goal: to concentrate users’ attention on only requirement iterations during the training, but make sure users can still apply their learning in overall prompt writing. Analyzing requirement defects in ROPE participants’ prompts, we observed a noticeable decrease in omission errors (from 5.6 to 3.2 per participant), but a slight increase in commission errors (from 0.5 to 0.7 per participant). This indicated that participants became more aware of requirements, but need additional training for requirements clarity and accuracy.
Figure 5. The pre-test and post-test overall scores between PE and ROPE conditions. (a) Overall scores for ROPE significantly improved from pre- to post-test; (b) ROPE achieved significantly higher learning gains (post -- pre) than PE (*** denotes $p<0.001$).
[…]
Bar chart comparing pre-test and post-test scores between PE (Prompt Engineering) and ROPE conditions. ROPE shows a statistically significant (p ¡ 0.001) improvement in overall scores from pre- to post-test, unlike PE.
@@
These prompts also revealed learning gain differences. A two-sample t-test showed that the ROPE achieved significantly higher learning gains compared to the PE group ($p<0.001$), as shown in Figure 5b. In fact, while ROPE significantly improved (19.1% as mentioned above), we did not observe much gain in PE group’s pre-to-post score (0.7%). The contrast suggested that novices indeed could not acquire requirement articulation skills through standard prompting training alone.